P4 (s) + 5O2 (g) P4O10 (g)
a) What mass of P4O10 forms when 2.7 g of O2 is used?
Step 2: Mass to moles
Step 3: Mole to Mole (Ratio)
Step 4: Mole to Mass
1
Expert's answer
2020-12-06T13:57:43-0500

P4 (s) + 5 O2 (g) = P4O10 (g)

M (O2) = 32 g/mol

M (P4O10) = 283.9 g/mol

n (O2) = 2.7 / 32 = 0.08 mol

n (P4O10) = 1/5 x n (O2) = 1/5 x 0.08 = 0.016 mol

m (P4O10) = 0.016 x 283.9 = 4.5 g